OpenStudy (anonymous):

Becky wants to make a sculpture in the shape of a rectangular prism for the science fair. The sculpture will be made of 16 21 cubic foot of clay and will have a base area of 4 5 square foot. How tall will the sculpture be? 20 21 foot 1 16 foot 1 21 foot 1 4 foot 4 21 foot

OpenStudy (kc_kennylau):

so, the volume is 16/21 cubic foot

OpenStudy (kc_kennylau):

and the base area is 4/5 square foot

OpenStudy (kc_kennylau):

we know that volume = base area x height

OpenStudy (kc_kennylau):

when you substitute those values in, you'd get 16/21 = 4/5 x height

OpenStudy (kc_kennylau):

I hope you can continue yourself :)

OpenStudy (anonymous):

so its what?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

so we can see that height = (16/21) / (4/5)

OpenStudy (kc_kennylau):

I hope you know how to do fraction division

OpenStudy (anonymous):

hahha i dont